For what value of b will 4x-8y=5 and 9y+bx=4 inconsistent?
Answer by Alan3354(69443)

For what value of b will 4x-8y=5 and 9y+bx=4 inconsistent?
-------------
4x-8y = 5
bx+9y = 4
-------------
Inconsistent --> parallel graphs --> equal slopes.
4x-8y = 5 --> m1 = 1/2
-----
bx+9y = 4
m2 = -b/9 = 1/2
b = -9/2